A box of blueberries is 2/5 full. Jane and friends had each eaten 1/10 of a box of blueberries. How many people ate blueberries?
This is not from a text book.

Answer by ptaylor(1894)


Let x=number of people who ate blueberries
I assume that when the friends started eating, the box was full (10/10)
Now we know the the number of people (x) multiplied by the amount they ate (1/10) of a box each subtracted from the full box (10/10) must equal the amount that's left (2/5). So our equation to solve is:
First lets change everything to decimals to make it easier:
10/10=1
1/10=0.1
2/5=0.4
1-0.1x=0.4 subtract 1 from both sides
-0.1x=-1+0.4
-0.1x=-0.6 Multiply both sides by -10
x=6-------------number that were eating on the blueberries
1-6(1/10)=2/5 or 4/10
10/10-6/10=4/10
4/10=4/10

Hope this helps---ptaylor